OpenStudy (liliakarina):

Distance bewtween two points is measured by the formula d= sqrt(X2-X1)^2 + (Y2-Y1)^2 I have this problem: 'What's the distance between A(3,4) and B(-3,-2) I did the operation and I get this result: 6.32 But the options are: 1.- 2 u ( u stands for units) 2.- 2 sqrt 6u 3.- 4u 4.- 6 sqrt 2u What am I doing wrong?

OpenStudy (mertsj):

\[\sqrt{(3-(-3))^2+(4-(-2))^2}=\sqrt{36+36}=\sqrt{72}=6\sqrt{2}\]

OpenStudy (mertsj):

It seems that they want you to leave the answer in exact radical form

OpenStudy (anonymous):

Your formula is missing a parentheses.\[d(P_1,P_2)=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\]

OpenStudy (liliakarina):

I see. So, P= point, i.e. A(3,4)?

OpenStudy (liliakarina):

@Mertsj So, I just get this answer with the radical form, I don't have to get the square root?

OpenStudy (anonymous):

A second indicator is that the x difference and the y difference are equal. This indicates a 45/45/90 triangle with sides of the ratio 1/1/sqrt2.

OpenStudy (mertsj):

Yes. Look at the list of answers. That will give you a clue.

OpenStudy (liliakarina):

Really helpful @Mertsj @AnimalAin !

OpenStudy (anonymous):

No sweat. Do math every day.